the following information for problems 38-40. The pharmacy stocks penicillin G vials with 5 million units in dry powder form. The reconstitution directions on the vial are as follows: Desired Concentration Solvent for Vial (units/ml) 250,000 500,000 750,000 1,000,000 (mL) 18.2 8.2 4.8 3.2 37 What volume of dry powder is contained in this container of penicillin G? If the vial was diluted with 18.2 mL of sterile water, what would the volume needed for 2 million units of penicillin G be? 38. If the vial was diluted with 4.8 mL, what would the volume needed for 2 million units of penicillin G be? 39 If the doctor orders KCl 40 mEq in 1 L of 0.9% NaCl over 8 hours, what is the flow rate in milliliters per hour? 40 The doctor orders 1 g of CaCl in 250 mL of DsW over 4 hr. How many drops per minute should be administered if the set is calibrated to deliver 40 gtt/mL? 41 Use the following information for problems 43 and 44 The doctor orders ampicillin 250 mg in 50 mL NS administered over 20 min every 6 hr using an IV administration set that delivers 15 gtt/mL. 42. What is the flow rate in drops per minute? 43 What is the flow rate in milliliters per minute?
Answer #1

37. Volume : 3.5gram (1g= 1.6million units)

38. 18.2ml= 250000,

3.64ml= 50000

3.64ml×4= 50000× 4

14.56ml= 200000units.

145.6ml= 2million units

39. 4.8ml= 750000

9.6ml= 1500000

3.2ml= 500000

12.8ml= 2million units.

40. formula = Amount of fluid × drop factors/ hours× minutes

= 1000 × 15/ 8× 60

= 31.25drops/ minutes

1ml= 20drops

1.5001ml = 31.25drops

1.5001ml.

41. = 250× 40/ 4× 60

= 41.66 drops / minutes.

42. = 50× 15/ 0.33(20min)× 60

= 37.87 drops.

43. 1ml= 20 drops

1.5ml= 30 drops

1.82ml= 37.87drops.